Write an equation of a line that passes through the points (2,3) and (2,-6).
Answer by Cromlix(4381)

Hi there,
Gradient = y2 - y1/x2 - x1
Gradient = -6 - 3/2 - 2 = -9/0
This is an undefined gradient.
Equation of line => x = 2
Hope this helps :-)